# Multiplicity dependence of (multi)strange hadrons in oxygen-oxygen collisions at $\sqrt{s_{\mathrm{NN}}}~{}=~{}7$ TeV using EPOS4 and AMPT

### v3

@@ III Results and Discussion
Figure 5 shows the $\phi/\pi$ and $\Xi/\pi$ ratios in the most central (0–5%) and peripheral (60–80%) $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~{}=~{}7$ TeV using EPOS4, AMPT-Def, and AMPT-SM models. The rise in $\phi/\pi$ and $\Xi/\pi$ ratios is more gradual, exhibiting a rapid increase at low $p_{\rm T}$ as compared to $K/\pi$ and $\Lambda/\pi$ ratios. Both ratios show weak centrality dependence at low $p_{\rm T}$ but strong centrality dependence at intermediate-$p_{\rm T}$ from all models. It is observed that at intermediate $p_{\rm T}$, the $\phi/\pi$ ratio is twice as large for AMPT-SM as compared to EPOS4 in the most central collisions. This may be because, at the intermediate-$p_{\rm T}$ region where the momentum of a particle becomes comparable or more than the mass of the strange quark, the production probability of $s\bar{s}$ pairs increases significantly. Additionally, $\phi$ mesons are not only produced from the fragmentation of excited strings in the initial collisions but also from the hadronic matter through various baryon-baryon, meson-baryon, and meson-meson scatterings. This leads to a corresponding rise in the observed $\phi/\pi$ ratio at intermediate $p_{\rm T}$. The difference in $\phi/\pi$ ratios in AMPT-SM and EPOS4 might also be attributed to the influence of the Lund string fragmentation parameters on the yields of $\phi$ mesons. AMPT-SM may have implemented these parameters more effectively, leading to a stronger preference for strange quark hadronization than in EPOS4 [66]. EPOS4 does not include a coalescence mechanism, and the $\phi/\pi$ ratio reflects the complex interplay between core-corona separation, hydrodynamics, collective flow, and particle production mechanisms across different $p_{\rm T}$ ranges. $\Xi/\pi$ ratios also show strong centrality dependence at the intermediate $p_{\rm T}$ in all models. Similar to the $\Lambda/\pi$ ratios, the $\Xi/\pi$ ratios exhibit a pronounced peak with a maximum of around 2–3 GeV/$c$ in most central collisions. A slightly higher prediction of $\Xi/\pi$ ratios in 0–5% central collisions in EPOS4 as compared to AMPT-SM may be due to the stronger flow effects in EPOS4.
@@
Figure 6 shows $p_{\rm T}$-differential $p/\pi$ and $\Lambda$/$\mathrm{K}^{0}_{\mathrm{S}}$ in the most central (0–5%) and peripheral (60–80%) $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~{}=~{}7$ TeV using all models. The $p_{\rm T}$-differential $p/\pi$ (the lightest baryon to the lightest meson) ratio serves as a proxy of the relative production of baryons compared to mesons. Both ratios show a strong centrality dependence at intermediate $p_{\rm T}$. A similar peak around 2–3 GeV/$c$ in most central collisions is observed in $p/\pi$ ratios from EPOS4 and AMPT-SM at intermediate $p_{\rm T}$.
@@
In AMPT, we observe a peak at intermediate $p_{\rm T}$ in central collisions, while a plateau appears at intermediate $p_{\rm T}$ in peripheral collisions similar to that observed in $\rm p+p$, $\rm p+Pb$, and $\rm Pb+Pb$ collisions at the LHC [8, 60, 61, 62, 63, 64]. The $\Lambda$/$\mathrm{K}^{0}_{\mathrm{S}}$ ratios also exhibit strong centrality dependence at intermediate $p_{\rm T}$.
@@
Figure 6: (Color online) $p_{\rm T}$-differential baryon-to-meson ($p/\pi$ and $\Lambda$/$\mathrm{K}^{0}_{\mathrm{S}}$) ratios in most central (0–5%) and the most peripheral (60–80%) $\rm O+O$ collisions using EPOS4 (top) and AMPT (bottom). Black is used for EPOS4, whereas red and green represent AMPT-Def and the AMPT-SM model, respectively. Solid and dashed bands represent the ratios at the most central (0–5%) and most peripheral (60–80%) centrality class, respectively.
@@
The enhancement in $\Lambda$/$\mathrm{K}^{0}_{\mathrm{S}}$ in 0–5% central collisions from EPOS4 is significantly larger than that predicted by the AMPT-SM model despite having a similar distribution. This discrepancy may be attributed to the strong radial flow implemented in EPOS4, which affects baryons like $\Lambda$ at higher multiplicities. In contrast, AMPT incorporates a certain degree of flow effects. This result aligns closely with the observed collision dynamics at this center-of-mass energy, indicating that the EPOS4 treatment of strangeness production mechanisms, especially for these particles, shows better consistency with the existing experimental data [37, 55].
@@ IV Conclusions
We present predictions of various observables for (multi)strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda$, $\Xi$, $\phi$, and $\Omega$) in $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~{}=~{}7$ TeV using the recently updated hydrodynamic-based EPOS4 and two different versions of the AMPT model. These models have been utilized to investigate the transverse momentum ($p_{\rm T}$) spectra, multiplicity dependence of the particle yield $\mathrm{d}N/\mathrm{d}y$, yield ratio of various strange hadrons, and $p_{\rm T}$-differential ratios of different strange meson and baryon relative to pions. EPOS4 performs well in predicting the enhancement of strangeness, while the current version of AMPT does not. None of the models can fully describe trends of strangeness enhancement observed in existing experimental data from other collision systems. A clear final-state multiplicity overlap is observed with $\rm p+p$, $\rm p+Pb$, and $\rm Pb+Pb$ collisions from all the models. The $p_{\rm T}$-differential strange hadron to pion ratios show a clear centrality dependence. AMPT predicts a stronger enhancement in the $\phi/\pi$ ratio compared to EPOS4 due to the effective implementation of the Lund string fragmentation parameters. The greater enhancement in the $\Lambda$/$\mathrm{K}^{0}_{\mathrm{S}}$ ratio in the intermediate-$p_{\rm T}$ region of EPOS4 compared to AMPT could be attributed to the implementation of complete hydrodynamic flow in EPOS4, which appears to play an important role and appears to be significantly more effective in reproducing existing experimental data from other collision systems, while AMPT incorporates some flow effects. The upcoming data on $\rm O+O$ collisions at the LHC will help constrain and refine the model parameters.
@@ V Acknowledgements
The authors would like to thank W. J. Llope and S.K. Grossberndt for careful reading and discussions. S. Kabana acknowledges partial support from the Agencia Nacional de Investigación y Desarrollo de Chile (ANID), Chile, with the grants ANID PIA/APOYO AFB230003, Chile, and ANID FONDECYT regular No 1230987, Chile. H. R. is supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant numbers ACI-1550300 OAC-2004571 (CSSI:X-SCAPE) and in part by the U.S. Department of Energy (DOE) under grant number DE-SC0021969.

### v4

@@ III Results and Discussion
Figure 5 shows the $\phi/\pi$ and $\Xi/\pi$ ratios in the most central (0–5%) and peripheral (60–80%) $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~=~7$ TeV using EPOS4, AMPT-Def, and AMPT-SM models. The rise in $\phi/\pi$ and $\Xi/\pi$ ratios is more gradual, exhibiting a rapid increase at low $p_{\rm T}$ as compared to $K/\pi$ and $\Lambda/\pi$ ratios. Both ratios show weak centrality dependence at low $p_{\rm T}$ but strong centrality dependence at intermediate-$p_{\rm T}$ from all models. It is observed that at intermediate $p_{\rm T}$, the $\phi/\pi$ ratio is twice as large for AMPT-SM as compared to EPOS4 in the most central collisions. This may be because, at the intermediate-$p_{\rm T}$ region where the momentum of a particle becomes comparable or more than the mass of the strange quark, the production probability of $s\bar{s}$ pairs increases significantly. Additionally, $\phi$ mesons are not only produced from the fragmentation of excited strings in the initial collisions but also from the hadronic matter through various baryon-baryon, meson-baryon, and meson-meson scatterings. This leads to a corresponding rise in the observed $\phi/\pi$ ratio at intermediate $p_{\rm T}$. The difference in $\phi/\pi$ ratios in AMPT-SM and EPOS4 might also be attributed to the influence of the Lund string fragmentation parameters on the yields of $\phi$ mesons. AMPT-SM may have implemented these parameters more effectively, leading to a stronger preference for strange quark hadronization than in EPOS4 [Tiwari:2020ult]. EPOS4 does not include a coalescence mechanism, and the $\phi/\pi$ ratio reflects the complex interplay between core-corona separation, hydrodynamics, collective flow, and particle production mechanisms across different $p_{\rm T}$ ranges. $\Xi/\pi$ ratios also show strong centrality dependence at the intermediate $p_{\rm T}$ in all models. Similar to the $\Lambda/\pi$ ratios, the $\Xi/\pi$ ratios exhibit a pronounced peak with a maximum of around 2–3 GeV/$c$ in most central collisions.
@@
Figure 6 shows $p_{\rm T}$-differential $p/\pi$ and $\Lambda$$\mathrm{K}^{0}_{\mathrm{S}}$ in the most central (0–5%) and peripheral (60–80%) $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~=~7$ TeV using all models. The $p_{\rm T}$-differential $p/\pi$ (the lightest baryon to the lightest meson) ratio serves as a proxy of the relative production of baryons compared to mesons. Both ratios show a strong centrality dependence at intermediate $p_{\rm T}$. A similar peak around 2–3 GeV/$c$ in most central collisions is observed in $p/\pi$ ratios from EPOS4 and AMPT-SM at intermediate $p_{\rm T}$.
@@
In AMPT, we observe a peak at intermediate $p_{\rm T}$ in central collisions, while a plateau appears at intermediate $p_{\rm T}$ in peripheral collisions similar to that observed in $\rm p+p$, $\rm p+Pb$, and $\rm Pb+Pb$ collisions at the LHC [5, 45, 49, 50, 51, 52]. The $\Lambda$$\mathrm{K}^{0}_{\mathrm{S}}$ ratios also exhibit strong centrality dependence at intermediate $p_{\rm T}$.
@@
Figure 6: (Color online) $p_{\rm T}$-differential baryon-to-meson ($p/\pi$ and $\Lambda$$\mathrm{K}^{0}_{\mathrm{S}}$) ratios in most central (0–5%) and the most peripheral (60–80%) $\rm O+O$ collisions using EPOS4 (top) and AMPT (bottom). Different colors are used to represent different models across various centrality classes.
@@
The enhancement in $\Lambda$$\mathrm{K}^{0}_{\mathrm{S}}$ in 0–5% central collisions from EPOS4 is significantly larger than that predicted by the AMPT-SM model despite having a similar distribution. This discrepancy may be attributed to the strong radial flow implemented in EPOS4, which affects baryons like $\Lambda$ at higher multiplicities. In contrast, AMPT incorporates a certain degree of flow effects. This result aligns closely with the observed collision dynamics at this center-of-mass energy, indicating that the EPOS4 treatment of strangeness production mechanisms, especially for these particles, shows better consistency with the existing experimental data [42, Werner:2013tya].
@@ IV Conclusions
We present predictions of various observables for (multi)strange hadrons ($\mathrm{K}^{0}_{\mathrm{S}}$, $\Lambda$, $\Xi$, $\phi$, and $\Omega$) in $\rm O+O$ collisions at $\sqrt{s_{\mathrm{NN}}}~=~7$ TeV using the recently updated hydrodynamic-based EPOS4 and two different versions of the AMPT model. These models have been utilized to investigate the transverse momentum ($p_{\rm T}$) spectra, multiplicity dependence of the particle yield $\mathrm{d}N/\mathrm{d}y$, yield ratio of various strange hadrons, and $p_{\rm T}$-differential ratios of different strange meson and baryon relative to pions. EPOS4 performs well in predicting the enhancement of strangeness, while the current version of AMPT does not. None of the models can fully describe trends of strangeness enhancement observed in existing experimental data from other collision systems. A clear final-state multiplicity overlap is observed with $\rm p+p$, $\rm p+Pb$, and $\rm Pb+Pb$ collisions from all the models. The $p_{\rm T}$-differential strange hadron to pion ratios show a clear centrality dependence. AMPT predicts a stronger enhancement in the $\phi/\pi$ ratio compared to EPOS4 due to the effective implementation of the Lund string fragmentation parameters. The stronger enhancement of the $\Lambda$$\mathrm{K}^{0}_{\mathrm{S}}$ ratio in the intermediate-$p_{\rm T}$ region in EPOS4 compared to AMPT can be attributed to the inclusion of a full hydrodynamic evolution in EPOS4, combined with its hadronization treatment (string fragmentation). Both factors enhance baryon production and lead to better agreement with experimental data from other collision systems. The upcoming data on $\rm O+O$ collisions at the LHC will help constrain and refine the model parameters.
@@ V Acknowledgements
The authors would like to thank W. J. Llope and S.K. Grossberndt for careful reading and discussions. H. R. is supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant numbers ACI-1550300 OAC-2004571 (CSSI:X-SCAPE) and in part by the U.S. Department of Energy (DOE) under grant number DE-SC0021969.